From nch@roe.ac.uk Fri Oct 15 11:58:29 2004 Date: Fri, 15 Oct 2004 11:41:22 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Harvey MacGillivray , Ian Bond , Mike Read , Nigel Hambly , Bob Mann Cc: CCs for WSA weekly meeting minutes distribution -- Clive Davenhall , Andrew Lawrence , Andy Adamson , Peter Shillan , John Taylor , Jim Emerson , Malcolm Stewart , Martin Hill , Mike Irwin , Peredur Williams Subject: WFAU WSA weekly meeting minutes: 15 October 2004 Minutes of WFCAM Science Archive meeting: 15th October 2004 ------------------------------------------------------------- ------------------------------------------------------------- Present: NCH, JMS, PMW, MAR, MCH, GPS, JDT, ETWS Apologies: JPE, AL, RGM, HMG DONM: 10am, Friday 22nd October, plate library Actions discharged: ------------------- ACTION: ETWS to add a TWiki note concerning UKLight high speed networking to the appropriate networking pages. Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to remember to action a team programmer to implement range checking in the low-level ingest codes. ETWS & NCH have been discussing this; likely that the new developer will be assigned this task to start them off. ACTION: JDT to liaise with MCH concerning an end-to-end test of the PAL deployment. SOAP services only appear to be not working; investigation is on-going; GPS volunteered to assist. ACTION: PMW to apply a little pressure to the premises team to solve the continuing problem of heat overload in C1. PMW reported bringing this up at the site services meeting this week; a quote has been obtained for additional airconditioning; action continues. ACTION: ALL to review risk register and think of any new internal or external risks that should be documented in the risk register. Defered until top level plan is explicitly defined. PMW tabled the VISTA risk register as an example. ACTION: JPE to design and set up centralised VDFS web pages. Progressing. (there is now a link to it from our WSA Twiki). - continuing; thanks for progressing these. Actions carried forward from 1/10/04 meeting: ----------------------------------------------- ACTION: HMG to oversee the removal of the DVD towers from C1. - CONTINUES (see Hardware below) Specific points and new actions: -------------------------------- Project management: JMS & PMW reported back from last week's VDMT meeting. The Vista URD release is imminent; in view of it's delay the schedule for reviews is being revisited and the mid-December milestone of a release of review documentation is defered. WSA progress was examined and found to be satisfactory. NCH and PMW noted that the software developer position is now filled; Iain MacCallum will be joining us on January the 17th. NCH reported that a WFCAM update from AA at JAC has been circulated around the UKIDSS consortium as follows: "Summary: barring one potentially major problem with the secondary, the work is on schedule still (and the camera now has a functional corrector plate). As you will understand, there remain some uncertainties and a further update will be sent as things develop. (i) The corrector plate was completed, delivered and installed into the cryostat on the 4th of October. This is great news since it will enable a much more complete first commissioning phase than would have been possible with the aperture plate, and will, other things being equal, considerably shorten the February/March engineering. (ii) The Zerodur fitting of one secondary-mirror support flexure sustained some damage at some point after the secondary was first mounted on the telescope. The cause of this is yet to be fully understood, and the long-term solution is not yet agreed. For the moment, the flexures have been rebonded back into the existing Zerodur rings (only one of which was significantly damaged) and the secondary has been installed back on the telescope. Because the piezo stacks deliver a significant kick to the flexures on powering up, the first week of commissioning will be done with power to the piezos switched off. Those parts of the on-sky commissioning / characterization which can be done without tip/tilt correction are being identified. (iii) The cause of the initial failure may have been a poor bond on one flexure, which allowed the flexure to disengage, thus putting an unacceptable stress on the one which in the end sustained damage. Although the current bonds have been load tested and appear good, there is a risk of further damage to the flexure mounts on powering up the tip/tilt stage. If so, we will complete as much commissioning as possible thereafter before removing WFCAM from the telescope and reworking the secondary. In this case the instrument would not be back on the telescope until February 2005. If not, then the first stage of commissioning will continue as planned - the instrument coming off the telescope in the week of the 15th of November for adjustment of the external tilt shims and autoguider lens, and then going back on through to the end of November. In this case, little commissioning work would remain for the Spring schedule slot, which would be used more for science observing. (iv) The cooldown begins this week, and WFCAM will be installed on the telescope next week. (v) First light is scheduled for the 21st of October !" Finally, NCH noted that the WFAU Revisable Grant renewal was due in mid-Novemeber and would contain bids for archive operators, as well as various other aspects of WFAU's programme. Work on this renewal case is likely to soak up the greater part of NCH, RGM and PMW effort over the next few weeks. Comments and issues arising from CASU fortnightly minutes: The team noted the minutes of the CASU meeting of 29/9/2004; in particular the summary "cunning plan" for WFCAM commissioning as follows: 21 Oct: First light; first data 3-4 days later 15 Nov: Completion of autoguider commissioning (inc. focus), followed by one week off to shim into focus 22 Nov: 8 days of on-sky tests followed by WFCAM off telescope until the beginning of February. Networking: ETWS reported appending some lines about the UKLight town meeting to the "Network for Non-Networkers" TWiki page. Hardware: NCH inquired as to the status of shifting the DVD jukeboxes since the new rack mount cabinet would be arriving soon; PMW agreed to follow this up. ACTION: PMW to chase up HMG to inquire as to the status of shifting the DVD jukeboxes. NCH reported that there is as yet no word from Eclipse concerning a date for the reconfigure and upgrade of the catalogue servers ACTION: NCH to chase up Eclipse to enquire as to status of the upgrade orders. Software: ETWS noted: "I've worked further on the GUI, the Archive and Programme History windows now have sortable columns. Also a new Programme Status Table showing the status of all CUs of all programmes colour coded is now available. And multiple CUs now can be written to a job list which can be scheduled at any time. The output which normally goes to stdout will be e-mailed to the operator." MAR noted that he has been working on the WSA poster for ADASS. NCH noted no progress on WSA software over the last week or so owing to attendance of the GAIA conference and other WFA related matters taking precedence. SSA: No further news this week. Astrogrid deployment: No progress on SOAP services on the archive network; MCH and GPS are continuing to investigate. There seems to be some security issue on the WSA private area network so MAR has enabled PSSA access on the grendel test system for MCH to investigate. Miscellaneous: Nothing else this week.